OpenStudy (anonymous):

for q.6 hw to find Isc

OpenStudy (anonymous):

isc=1.5 for ur values

OpenStudy (anonymous):

thanx..............

OpenStudy (anonymous):

@Tejaswee: by short ckting output terminal

OpenStudy (anonymous):

what are your values...

OpenStudy (nurali):

Question 6 Formula Vout = Vin*(R3-B*R1*R3)/(R3-(B*R1*R3)+R1+R2+A*R2) i1=V0/(R3-B*R1*R3) and Vth is same as Vout. Isc=vIN*(1-B*R1)/(A*R2+R1+R2) Rth=VO/iSC where VO is Vth

OpenStudy (nurali):

If correct please click best response